OpenStudy (anonymous):

Is (0, 2) a solution to y = -x + 2?

OpenStudy (dcolley):

yes

OpenStudy (anonymous):

thanks dcolley

OpenStudy (dcolley):

your welcome

OpenStudy (anonymous):

y=-x+2 0=-2+2 0=0

OpenStudy (dcolley):

0 is the value

OpenStudy (anonymous):

this implies that 0,2 is the solution of given system

OpenStudy (dcolley):

i meant 0 is the x value so is it still right cause i just plugged in the values

OpenStudy (anonymous):

ofcourse 0 the value of x and 2 is the value of y